# Large Language Model LRM Large Reasoning Model

### v2

@@ Abstract
Large Language Model (LLM)-based agents solve complex tasks through iterative reasoning, exploration, and tool-use, a process that can result in long, expensive context histories. While state-of-the-art Software Engineering (SE) agents like OpenHands or Cursor use LLM-based summarization to tackle this issue, it is unclear whether the increased complexity offers tangible performance benefits compared to simply omitting older observations. We present a systematic comparison of these strategies within SWE-agent on SWE-bench Verified across five diverse model configurations. We find that a simple observation-masking strategy halves cost relative to a raw agent while matching, and sometimes slightly exceeding, the solve rate of LLM summarization. For example, with Qwen3-Coder 480B, masking improves solve rate from 53.8% (raw agent) to 54.8%, while remaining competitive with summarization at a lower cost. These results suggest that, at least within SWE-agent on SWE-bench Verified, the most effective and efficient context management can be the simplest. We release code and data for reproducibility. [^1] Data: https://huggingface.co/datasets/the-complexity-trap-19fna01lf1Y/main-experiments [^2] Code: https://anonymous.4open.science/r/complexity-trap-6167/
@@ 1 Introduction
The ambition to create autonomous agents that can independently handle complex SE tasks is rapidly becoming a reality. These agents, powered by LLMs, typically operate in an iterative loop Yao et al. (2023); Wang et al. (2024), at each turn they reason about the current state, devise a plan, and execute a tool (e.g., read a file, run tests). The output, or observation, from this tool is then added to the agent’s context for the next turn, extending its problem-solving trajectory (Figure˜3). This agentic loop acts as a powerful test-time scaling mechanism Snell et al. (2025); Liu et al. ; Wu et al. (2024), utilizing the reasoning capabilities Wei et al. (2022) of LLMs at each turn while grounding them through environment responses.
However, this iterative context expansion presents a fundamental tradeoff between cost and capability, or effectiveness and efficiency. As the agent’s trajectory grows, calls to the LLM become prohibitively expensive due to token-based pricing, and inefficient due to the quadratic attention complexity in the widespread Transformer architecture Vaswani et al. (2017). More critically, even with context windows exceeding 1M tokens, LLMs suffer from the "lost in the middle" problem Liu et al. (2024). While LLMs can process large context windows, they cannot properly make use of relevant information buried within their vast context Modarressi et al. ; Hong et al. (2025). This challenge is acutely amplified in the SE domain, where tool observations are notoriously verbose and noisy Lindenbauer et al. (2025). A single action can yield thousands of tokens, whether from reading an entire source file, running a recursive directory listing, or a lengthy test suite log. Concretely, observation tokens make up around $84\%$ of an average SWE-agent turn Yang et al. (2024) (Figure˜1) in our preliminary experiments (Section˜D.4) on SWE-bench Lite-50 Jimenez et al. (2024); Badertdinov et al. (2024). Due to this, targeting environment observations explicitly provides a strong baseline for LLM-agent context management.
@@
A natural baseline approach directly targets verbose environment observations through observation masking strategies that explicitly omit tool outputs. State-of-the-art systems like SWE-agent Yang et al. (2024) and SWE-Search Antoniades et al. (2025) have adopted such relatively simple approaches. In parallel, an increasingly popular and more sophisticated solution is to prompt an LLM to perform trajectory summarization, replacing parts of the agent’s history with condensed summaries (Figure˜3). This approach is adopted by prominent open-source and proprietary SE agents like OpenHands Wang et al. (2025) and Cursor Inc. . Open-source implementations of both the observation masking and LLM summarization approaches rely on heuristic triggers such as fixed context window sizes or turn thresholds. Thus, the key difference is whether they discard or condense old context. Despite the critical impact of this choice on agent cost and performance, the relative trade-offs between these approaches remain largely unexplored.
@@
In this work, we present a systematic comparison focused on the efficiency of context management strategies within SWE-agent Yang et al. (2024). For this, we analyze the performance of LLM-based trajectory summarization and a simple observation masking baseline for context management with respect to efficiency (cost in USD) and effectiveness (solve rate on the challenging and industry-standard SWE-bench Verified Chowdhury et al. (2024) benchmark). We implement an LLM-based summarization in SWE-agent Yang et al. (2024) and utilize the OpenHands-style prompt (Figure˜11) to enable a controlled, within-scaffold comparison. We evaluate these strategies within the SWE-agent Yang et al. (2024) scaffold across a diverse spectrum of experimental conditions, including model families, model sizes, and reasoning regimes.
We find that the simple observation-masking approach more than halves the cost compared to the raw agent while matching and sometimes exceeding the performance of the LLM-based summarization and the raw agent within SWE-agent. A finding that challenges current trends in the field and demonstrates that in the quest for more capable LLM agents, complexity is not always the answer, sometimes the simplest solution is also the most effective.
@@
Figure 2: The effectiveness versus efficiency tradeoff for context management strategies within SWE-agent Yang et al. (2024) on SWE-bench Verified Chowdhury et al. (2024). The plot compares solve rate (y-axis, $\uparrow$) against the average cost per trajectory (x-axis, $\downarrow$) for different model configurations. We test each configuration with three strategies: raw agent (baseline, ), LLM-Summary (), and observation masking (). Across all models, the observation-masking strategy consistently occupies the most efficient frontier, achieving solve rates competitive with, and sometimes superior to, the LLM-Summary strategy. With Qwen3-Coder 480B Team (2025); Hui et al. (2024), the best-performing model in our experiments, Observation Masking is not only 52% cheaper than the Raw Agent baseline but also improves on the solve rate by 2.6 %. Moreover, it even reduces the cost per instance compared to LLM-Summary by \$0.03 (\$15 across 500 instances).
@@ 2 Related Work
Current SE agent research mostly focuses on improving the effectiveness of SE agents by scaling training data Jain et al. (2025); Pan et al. (2024); Yang et al. (2025), selecting the most promising of multiple attempts Pan et al. (2024); Jain et al. (2025); Zhang et al. (2025); Antoniades et al. (2025); Aggarwal et al. (2025); Xia et al. , providing execution-free or execution-based feedback to the agent through critics Antoniades et al. (2025); Pan et al. (2024); Jain et al. (2025); Zhang et al. (2025); Shinn et al. (2023), or enhancing the agent’s planning capabilities through explicit search strategies Antoniades et al. (2025); Aggarwal et al. (2025). While these methods improve solve rate, they come at the cost of increased inference costs and thus reduced efficiency. Efficient context management for SE agents, on the other hand, has thus far received little attention. In this work, we investigate whether complex summarization strategies are necessary for efficient context management in SE agents. For this we experiment with SWE-agent Yang et al. (2024). Moreover, we show that without any such context management the instance cost routinely more than doubles, positioning research on efficient agents as pivotal for an economically feasible, large-scale deployment of SE agents.
Recently, several works have conducted in-depth analyses on the effect of the context size on the LLMs performance Liu et al. (2024); Hong et al. (2025); Modarressi et al. . These consistently show that LLMs increasingly struggle to effectively utilize the context provided with increased context size. Despite the critical importance of context management for agent performance and deployment costs, existing work treats it as an implementation detail rather than a first-class research question. A recent exception, MEM1 Zhou et al. , explores dynamic state management for multi-hop QA Yang et al. ; Kwiatkowski et al. (2019); Jin et al. (2025) and web navigation Yao et al. tasks. However, the authors do not compare to omission-based approaches. Furthermore, the benchmarks used result in relatively short trajectories (hundreds of tokens) Zhou et al. compared to SE agent trajectories that routinely are orders of magnitude larger Lindenbauer et al. (2025).
@@ 3 Experimental Configuration
Figure 3: Overview of the context management strategies evaluated in our work. Box heights indicate the number of tokens in that portion of a typical trajectory. (a) The baseline ReAct-style Yao et al. (2023) trajectory, where the context grows with each action-observation pair. (b) LLM-based summarization condenses older turns into a running summary and preserving a few recent turns in full (e.g., OpenHands Wang et al. (2025)). (c) Observation masking replaces observations older than a fixed window of M turns (here, M=1) with a placeholder (e.g., SWE-agent Yang et al. (2024)).
In our experiments we cover one popular, representative approach for observation omission and LLM-based summarization: (1) environment observation masking through a rolling window Yang et al. (2024); Antoniades et al. (2025), and (2) prompt-based LLM summarization Wang et al. (2025); Inc. . In the following, we will define our chosen approaches in detail.
@@ 3.1 Context Management Strategies
Raw baseline agent.
The agent scaffolds we investigate use either ReAct Yao et al. (2023) or CodeAct Wang et al. (2024) as an agent framework. In these frameworks, the agent’s history, or trajectory, is a sequence of interactions with an environment. We formalize the trajectory at the end of turn t1t-1, denoted $\tau_{t-1}$, as a sequence of tokens:
@@
Here, $p_{i}$ is a placeholder text representing the masked observation, such as “Previous 8 lines omitted for brevity.”. In the following turns, the agent LLM then conditions on this condensed history $\tau^{\prime}_{t-1}$ to produce $(r_{t},a_{t})$. This approach, following SWE-agent Yang et al. (2024), retains the complete reasoning chain while reducing distant observation fidelity. While this strategy reduces the speed at which the tokens in $\tau$ grow, it does not solve the issue of indefinite growth.
LLM Summaries.
@@ 3.2 Experimental Configuration
We conduct a rigorous within-scaffold study using SWE-agent Yang et al. (2024), covering distinct (1) model families, (2) model sizes, (3) model licenses, (4) reasoning regimes (thinking and non-thinking). Concretely, we use Qwen3-32B [^3] https://huggingface.co/Qwen/Qwen3-32B Team (2025) in thinking and non-thinking mode with a context window of 122122K tokens using YaRN Peng et al. , Qwen3-Coder-480B-A35B-Instruct-FP8444https://huggingface.co/Qwen/Qwen3-Coder-480B-A35B-Instruct-FP8 Team (2025); Hui et al. (2024), and Gemini 2.5 Flash [^5] API Version: gemini-2.5-flash Comanici et al. (2025) in thinking and non-thinking mode. We conduct all experiments with on SWE-agent Yang et al. (2024) on SWE-bench Verified Chowdhury et al. (2024) unless otherwise specified. We run all our experiments on a shared cluster of eight NVIDIA H200 GPUs, each equipped with 141 GB of HBM, and a total of 8 TB local disk storage. For the Qwen-32B Team (2025) models we use two H200 GPUs, and 15 SWE-agent Yang et al. (2024) inference workers. We choose a conservative number of workers, because we may encounter long trajectories with context sizes >100>100K tokens and must account for this in our KV-cache estimates. For Qwen3-Coder 480B Team (2025); Hui et al. (2024), we use all eight GPUs and 35 inference workers. We use vLLM Kwon et al. (2023) to serve the Qwen models on our cluster. In our experiments with Gemini we use eight inference workers on SWE-agent Yang et al. (2024) due to quota restrictions. For further details see Appendix˜A.
Our preliminary experiments (Section˜D.4) indicate that the number of turns in the trajectory may influence the behavior of the context management strategies we investigate. Due to this, we experiment with long trajectories and thus set the turn limit to 250 in our main experiments. For the Observation Masking strategy we use a rolling window size of $M=10$ in our main experiments, because it resulted in the best performance with SWE-agent Yang et al. (2024) in our experiments (Section˜D.1). The LLM-Summary strategy we implement in SWE-agent Yang et al. (2024) uses a slightly modified version of the OpenHands-style prompt (Appendix˜B). In contrast to OpenHands Wang et al. (2025), we summarize 21 turns at once $N=21$ and retain only the last ten $M=10$ turns. Besides aligning the number of unaltered tail turns for the Observation Masking and LLM-Summary strategies, we also found that $M=10$ offered the best performance for the LLM-Summary strategy in our experiments (Section˜D.2). For the agent model we use a temperature of 0.8, and for the summary model a temperature of zero. In contrast to experiments with Qwen3-32B Team (2025) thinking, we restrict the thinking budget of Gemini 2.5 Flash to zero or 800 tokens (denoted as thinking) due to cost constraints.
@@ 4 Main Results
Our main experiments within SWE-agent Yang et al. (2024) evaluate three context management strategies, with results summarized in Table 1. The results reveal a central finding that holds robustly across the diverse conditions we tested, spanning model families (Qwen Team (2025); Hui et al. (2024), Gemini Comanici et al. (2025)), sizes (32B, 480B), licenses (open-source, proprietary), and reasoning regimes (with and without thinking). The context management strategies evaluated significantly reduce costs, without significantly reducing solve rate performance. For further details see Appendix˜C.
@@ 4.1 The Universal Benefit of Context Management
Our first and most foundational finding, reinforcing the motivation of this study, is that context management is not merely an optimization but a necessity. As shown in Table˜1, leaving the agent’s context to grow unchecked (the Raw Agent baseline) is consistently the most expensive strategy. In all five of our experimental configurations, both Observation Masking and LLM-Summary drastically reduce the cost per instance, in most cases by more than 50%.
Furthermore, this efficiency does not necessarily come at the cost of performance. In three of our five setups, the most efficient strategy also achieved a higher solve rate than the Raw Agent baseline. This demonstrates that beyond a certain point, more context becomes a liability rather than an asset, aligning with the "lost in the middle" problem Liu et al. (2024). Therefore it is critical to question which approach offers the best trade-off between effectiveness and efficiency.
Table 1: Comparison of context management strategies with 95% bootstrap confidence intervals. We report change and significance (†) compared to the Raw Agent. We report Solve Rate (effectiveness, $\uparrow$) and Instance Cost (efficiency, $\downarrow$). For each model, we boldface the best-performing context management strategy for each metric. All experiments use SWE-agent Yang et al. (2024) on SWE-bench Verified Chowdhury et al. (2024). Further details in Appendix˜C.

### v3

# Deep Learning for Code in the Agentic Era

@@ Abstract
Large Language Model (LLM)-based agents solve complex tasks through iterative reasoning, exploration, and tool-use, a process that can result in long, expensive context histories. While state-of-the-art Software Engineering (SE) agents like OpenHands or Cursor use LLM-based summarization to tackle this issue, it is unclear whether the increased complexity offers tangible performance benefits compared to simply omitting older observations. We present a systematic comparison of these approaches with SWE-agent on SWE-bench Verified across five diverse model configurations. Moreover, we show initial evidence of our findings generalizing to the OpenHands agent scaffold. We find that a simple environment Observation Masking strategy halves cost relative to the Raw Agent while matching, and sometimes slightly exceeding, the solve rate of LLM-Summary. Additionally, we introduce a novel hybrid approach that further reduces costs by 7% and 11% compared to just Observation Masking or LLM-Summary, respectively. Our findings raise concerns regarding the trend towards pure LLM-Summary and provide initial evidence of untapped cost reductions by pushing the efficiency-effectiveness frontier. We release code and data for reproducibility. [^1] Data: https://huggingface.co/datasets/JetBrains-Research/the-complexity-trap [^2] Code: https://github.com/JetBrains-Research/the-complexity-trap
@@ 1 Introduction
The ambition to create autonomous agents that can independently handle complex SE tasks is rapidly becoming a reality. These agents, powered by LLMs, typically operate in an iterative loop [39, 29], at each turn they reason about the current state, devise a plan, and execute a tool (e.g., read a file, run tests). The output, or observation, from this tool is then added to the agent’s context for the next turn, extending its problem-solving trajectory (Figure˜3). This agentic loop acts as a powerful test-time scaling mechanism [25, 17, 32], utilizing the reasoning capabilities [31] of LLMs at each turn while grounding them through environment responses.
However, this iterative context expansion presents a fundamental tradeoff between cost and capability, or effectiveness and efficiency. As the agent’s trajectory grows, calls to the LLM become prohibitively expensive due to token-based pricing, and inefficient due to the quadratic attention complexity in the wide-spread Transformer architecture [28]. More critically, even with context windows exceeding 1M tokens, LLMs suffer from the "lost in the middle" problem [16]. While LLMs can process large context windows, they cannot properly make use of relevant information buried within their vast context [19, 7]. This challenge is acutely amplified in the SE domain, where tool observations are notoriously verbose and noisy [15]. A single action can yield thousands of tokens, whether from reading an entire source file, running a recursive directory listing, or a lengthy test suite log. Concretely, observation tokens make up around $84\%$ of an average SWE-agent turn [35] (Figure˜1) in our preliminary experiments (Section˜D.4) on SWE-bench Lite-50 [11, 3]. Due to this, targeting environment observations explicitly provides a strong baseline for LLM-agent context management.
@@
A natural baseline approach directly targets verbose environment observations through observation masking strategies that explicitly omit tool outputs. State-of-the-art systems like SWE-agent [35] and SWE-Search [2] have adopted such relatively simple approaches. In parallel, an increasingly popular and more sophisticated solution is to prompt an LLM to perform trajectory summarization, replacing parts of the agent’s history with condensed summaries (Figure˜3). This approach is adopted by prominent open-source and proprietary SE agents like OpenHands [30] or Cursor [9]. Open-source implementations of both the observation masking and LLM summarization approaches rely on heuristic triggers such as fixed context window sizes or turn thresholds. Thus, the key difference is whether they discard or condense old context. Despite the critical impact of this choice on agent cost and performance, the relative trade-offs between these approaches remain largely unexplored.
@@
In this work, we present a systematic comparison focused on the efficiency of context management strategies. For this, we analyze the performance of representative open-source Observation Masking and LLM-Summary implementations with respect to efficiency (cost in USD) and effectiveness (solve rate on the challenging and industry-standard SWE-bench Verified [5] benchmark). We capitalize these names throughout to denote the specific strategies formally defined in Section˜3.1. To enable controlled experimentation across model configurations, we implement LLM-Summary in SWE-agent [35] and adapt the OpenHands’ LLM-Summary prompt (Figure˜11). We evaluate these strategies within the SWE-agent [35] scaffold and probe for generality on OpenHands [30]. Our experiments span model families, model sizes, licenses (open-weights vs. proprietary), and reasoning regimes (thinking vs. non-thinking).
We find that both Observation Masking and LLM-Summary more than halve the cost compared to the Raw Agent. Furthermore, LLM-Summary is unable to consistently or significantly outperform the simple Observation Masking strategy across all model configurations. We show initial evidence of these findings generalizing to OpenHands. Furthermore, we introduce a novel hybrid approach that further reduces costs by 7% and 11% compared to just Observation Masking or LLM-Summary, respectively. These findings challenge current trends toward pure LLM-Summary and demonstrate that pure LLM-Summary strategies likely leave considerable cost savings untapped.
@@
Figure 2: The effectiveness versus efficiency tradeoff for context management strategies within SWE-agent [35] on SWE-bench Verified [5]. The plot compares solve rate (y-axis, $\uparrow$) against the average cost per trajectory (x-axis, $\downarrow$) for different model configurations. We test each configuration with three strategies: Raw Agent (baseline, ), LLM-Summary (), and Observation Masking (). Across all models, the Observation Masking strategy consistently occupies the most efficient frontier, achieving solve rates competitive with, and sometimes superior to, the LLM-Summary strategy. With Qwen3-Coder 480B [27, 8], the best-performing model in our experiments, Observation Masking is not only 52% cheaper than the Raw Agent baseline but also improves on the solve rate by 2.6 %. Moreover, it even reduces the cost per instance compared to LLM-Summary by \$0.03 (\$15 across 500 instances).
@@ 2 Related Work
Current SE agent research mostly focuses on improving the effectiveness of SE agents by scaling training data [10, 22, 36], selecting the most promising of multiple attempts [22, 10, 40, 2, 1, 33], providing execution-free or execution-based feedback to the agent through critics [2, 22, 10, 40, 24], or enhancing the agent’s planning capabilities through explicit search strategies [2, 1]. While these methods improve solve rate, they come at the cost of increased inference costs and thus reduced efficiency. Efficient context management for SE agents, on the other hand, has thus far received little attention. In this work, we investigate whether complex summarization strategies are necessary for efficient context management in SE agents. For this we experiment with SWE-agent [35] and OpenHands [30].
Recently, several works have conducted in-depth analyses on the effect of the context size on the LLMs performance [16, 7, 19]. These consistently show that LLMs increasingly struggle to effectively utilize the context provided with increased context size. Despite the critical importance of context management for agent performance and deployment costs, existing work treats it as an implementation detail rather than a first-class research question. A recent exception, MEM1 [41], explores dynamic state management for multi-hop QA [37, 13, 12] and web navigation [38] tasks. However, the authors do not compare to omission-based approaches. Furthermore, the benchmarks used result in relatively short trajectories (hundreds of tokens) [41] compared to SE agent trajectories that routinely are orders of magnitude larger [15].
Concurrently with our work, Xiao et al. [34] propose an LLM-Summary variant for efficient SE agent context management. However, they do not compare with an Observation Masking baseline. Most closely related, their “Delete” baseline is an LLM-Summary variant that deletes full turns instead of summarizing them. This baseline does not take advantage of SE trajectories skewing towards environment observations (Figure˜1) and breaks the agent reasoning trace across turns, degrading agent downstream performance. Irrespective of this, their results show that this baseline is more efficient than LLM-Summary at a comparable downstream performance. Additionally, Lu et al. [18] investigate using LLM-Summary for tackling context window limitations in RL training of SE and Computer-Use Agent (CUA). Tang et al. [26] on the other hand, achieve impressive performance with Observation Masking for RL training and inference of deep research and CUA agents. This underscores the timeliness of our research and that our findings likely generalize to deep research and CUA.
@@ 3 Experimental Configuration
Figure 3: Overview of the context management strategies evaluated in our work. Box heights indicate the number of tokens in that portion of a typical trajectory. (a) The baseline ReAct-style [39] trajectory, where the context grows with each action-observation pair. (b) LLM-based summarization condenses older turns into a running summary and preserving a few recent turns in full (e.g., OpenHands [30]). (c) Observation masking replaces observations older than a fixed window of M turns (here, M=1) with a placeholder (e.g., SWE-agent [35]).
In our experiments we investigate popular approaches to observation omission and LLM-based summarization through representative open-source implementations: (1) environment Observation Masking through a rolling window [35, 2], and (2) prompt-based LLM-Summary [30, 9]. In the following, we will define our chosen approaches in detail.
@@ 3.1 Context Management Strategies
Raw Agent.
The agent scaffolds we investigate use either ReAct [39] or CodeAct [29] as an agent framework. In these frameworks, the agent’s history, or trajectory, is a sequence of interactions with an environment. We formalize the trajectory at the end of turn t1t-1, denoted $\tau_{t-1}$, as a sequence of tokens:
@@
Here, $p_{i}$ is a placeholder text representing the masked observation, such as “Previous 8 lines omitted for brevity.”. In the following turns, the agent LLM then conditions on this condensed history $\tau^{\prime}_{t-1}$ to produce $(r_{t},a_{t})$. This approach, following SWE-agent [35], retains the complete reasoning chain while reducing distant observation fidelity. While this strategy reduces the speed at which the tokens in $\tau$ grow, it does not solve the issue of indefinite growth.
LLM-Summary.
@@ 3.2 Experimental Configuration
We conduct a rigorous, comparative study focusing on SWE-agent [35] and probing for generality with OpenHands [30] (v0.43.0). Our experiments cover diverse configurations spanning (1) model families, (2) model sizes, (3) model licenses (open-weights vs. proprietary), (4) reasoning regimes (thinking and non-thinking). Concretely, we use Qwen3-32B [27] [^3] https://huggingface.co/Qwen/Qwen3-32B in thinking and non-thinking mode with a context window of 122K tokens using YaRN [23], Qwen3-Coder-480B-A35B-Instruct-FP8 [27, 8] [^4] https://huggingface.co/Qwen/Qwen3-Coder-480B-A35B-Instruct-FP8 with its default context window of 256K tokens, and Gemini 2.5 Flash [6] [^5] API Version: gemini-2.5-flash with its default context window of 1M tokens in thinking and non-thinking mode. We conduct all experiments on SWE-bench Verified [5] unless otherwise specified. We run all our experiments on a shared cluster of eight NVIDIA H200 GPUs, each equipped with 141 GB of HBM, and a total of 8 TB local disk storage. For the Qwen-32B [27] models we use two H200 GPUs, and 15 SWE-agent [35] inference workers. We choose a conservative number of workers, because we may encounter long trajectories with context sizes >100>100K tokens and must account for this in our KV-cache estimates. For Qwen3-Coder 480B [27, 8], we use all eight GPUs and 35 inference workers. We use vLLM [14] to serve the Qwen models on our cluster. In our experiments with Gemini we use eight inference workers on SWE-agent [35] and five on OpenHands [30] due to quota restrictions. For further details see Appendix˜A.
Our preliminary experiments (Section˜D.4) indicate that the number of turns in the trajectory may influence the behavior of the context management strategies we investigate. Due to this, we experiment with long trajectories and thus set the turn limit to 250 in our experiments unless otherwise specified. For the Observation Masking strategy we use a rolling window size of $M=10$ in our main experiments, because it resulted in the best performance with SWE-agent [35] in our experiments (Section˜D.1). The LLM-Summary strategy we implement in SWE-agent [35] uses a slightly modified version of the OpenHands-style prompt (Appendix˜E). In contrast to OpenHands’ baseline configuration [30], we summarize 21 turns at once ($N=21$) and retain only the last ten ($M=10$) turns. Besides aligning the number of unaltered tail turns for the Observation Masking and LLM-Summary strategies, we also found that $M=10$ offered the best performance for the LLM-Summary strategy in our experiments (Section˜D.2). For the agent model we use a temperature of 0.8, and for the summary model a temperature of zero. In contrast to experiments with Qwen3-32B [27] thinking, we restrict the thinking budget of Gemini 2.5 Flash to zero or 800 tokens (denoted as thinking) due to cost constraints.
@@ 4 Main Results
Our main experiments within SWE-agent [35] evaluate three context management strategies, with results summarized in Table 1. The results reveal two central findings that hold robustly across the diverse conditions we tested. First, both Observation Masking and LLM-Summary significantly reduce costs, without significantly reducing solve rate performance. Second, LLM-Summary does not consistently, or significantly outperform Observation Masking on efficiency or effectiveness. For further details see Appendix˜C.
@@ 4.1 The Universal Benefit of Context Management
Our first and most foundational finding, reinforcing the motivation of this study, is that context management is not merely an optimization but a necessity. As shown in Table˜1, leaving the agent’s context to grow unchecked (the Raw Agent baseline) is consistently the most expensive strategy. In all experimental configurations where trajectories are long enough to benefit from efficient context management, both Observation Masking and LLM-Summary significantly reduce the cost per instance, in most cases by more than 50%. We discuss the outlying behavior of Qwen3-32B (thinking) in Appendix˜B.
Furthermore, this efficiency does not necessarily come at the cost of performance. In three of our five setups, the most efficient strategy also achieved a higher solve rate than the Raw Agent baseline. This demonstrates that beyond a certain point, more context becomes a liability rather than an asset, aligning with the "lost in the middle" problem [16]. Therefore it is critical to question which approach offers the best trade-off between effectiveness and efficiency.
Table 1: Comparison of context management strategies with 95% bootstrap confidence intervals. We report change and significance (†) compared to the Raw Agent. We report Solve Rate (effectiveness, $\uparrow$) and Instance Cost (efficiency, $\downarrow$). For each model, we boldface the best-performing context management strategy for each metric. All experiments use SWE-agent [35] on SWE-bench Verified [5]. Further details in Appendix˜C.